# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ Appendix I Machine Learning Efficiency Models
These days, inspired from diffusion models in the image and video domains, much work relies on the idea of latent diffusion. Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space differs from that.
@@
Latent diffusion models first encode the data and map it into a latent space. The diffusion model itself is then trained in that latent space. Hence, the performance of the diffusion model directly depends on a second, separate model, with a separate training procedure. TabSyn uses a VAE model to encode mixed-type data into a common continuous space that is not lower-dimensional, so as to minimize reconstruction errors. Any reconstruction errors caused by the incapability of the VAE in turn reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model. This suggests that we would want to train a high capable encoder/decoder, which adds additional training costs. Figure 3 shows that latent diffusion is not necessarily more efficient in the tabular data domain. In particular, if the latent space is not lower-dimensional to minimize reconstruction error, then sampling speed is not improved.
@@
We further hypothesize that much tabular data, due to the lack of redundancy and spatial or sequential correlation, is difficult to summarize efficiently in a joint latent space. Hence, compared to other domains, larger VAEs and higher-dimensional latent spaces are required, increasing the training time. Also, there is the risk of the VAE not picking up on subtle correlations within the data or distorting existing correlations by mapping into the latent space. Any correlations not properly encoded in the latent space, cannot be learned or exploited by the diffusion model. Since we optimize the VAE on an average loss, its reconstruction and encoding performance of, for instance, minority classes or extreme values in long-tailed distributions is likely lacking. This makes the job of the diffusion model more difficult, if not impossible.
@@
Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix A and B). This is a crucial part of modeling mixed-type data. Using a VAE to define a diffusion process in latent space only shifts the necessity for homogenization to the VAE training process. Not balancing different feature- or data-types and their losses induces an implicit importance weight for each feature. Thus, the VAE may sacrifice the reconstruction quality of some features in favor of others (Kendall et al., 2018; Ma et al., 2020).
@@
To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine the worst feature-specific sample quality and other metrics that directly benefit from the model generating all features well. Our results in Table 9 show that, latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix G). In particular, the attained maximum metrics indicate that TabSyn has issues modeling all features and their correlations sufficiently well. This supports our argument that a homogenization of data types is of crucial importance to avoid having the model implicitly favor one feature over another.
@@
Table 9: A comparison of the CDTD model to latent diffusion (TabSyn). We average each metric over five sampling seeds and as a robust measure report the median over the ablation datasets acsincome, adult, beijing and churn. Abs. diff. in corr. matrices refers to the absolute differences in the correlation matrices between ground truth and synthetic data. The maximum, minimum and mean are taken across features.
Detection score L2 dist. of corr. JSD WD Abs. diff. in corr. matrices
@@
TabSyn 0.772 0.479 0.005 0.018 0.046 0.003 0.006 0.017 0 0.133
CDTD (per type) 0.566 0.131 0.001 0.012 0.022 0.001 0.003 0.007 0 0.052
improvement over TabSyn 1.364 3.656 5.000 1.500 2.091 3.000 2.000 2.429 0 2.558
Appendix N Comparison to Related Work
Table 10 summarizes our comparison of CDTD to the selected diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi. Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix M). TabSyn is the only other model besides CDTD that avoids one-hot encoding categorical features by using embeddings. This improves the scalability to a higher number of categories without blowing up the input dimensions. Although both models utilize embeddings, TabSyn’s generative capabilities are more constrained by jointly encoding all features in a latent space. As such, it is still less flexible than CDTD, in particular when modeling very unbalanced categorical data. Information on rare categories may easily be cut off in favor of attributing more capacity in the latent space to more prominent categories or features. It should also be noted that TabSyn is the only model that makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
The CDTD model is the first to utilize adaptive and type- or feature-specific noise schedules to model tabular data. Further, we take great care in homogenizing categorical and continuous features throughout the training process, including the model initialization (see Appendix A and B). No other model attempts balancing the different features types. This is problematic as it suggests that other models may suffer from feature-specific induced implicit importance weights that impact both training and generation processes. Hence, the sample quality of some features may be unintentionally sacrificed in favor of increasing the sample quality of other features (Kendall et al., 2018; Ma et al., 2020). Note that this also applies to TabSyn: Even though their diffusion model avoids this issue by relying on a single type of loss due to the continuous latent space, the VAE training process does not account for any balancing issues between the two data types. Hence, the balancing issue is not eliminated but got only shifted to the encoder VAE.
@@
defined in feature space avoids one-hot encoding balances feature types adaptive noise schedule type- or feature- specific noise schedules diffusion in continuous space
CoDi
TabDDPM
TabSyn $\ast$
CDTD (ours)
Appendix O Examples of Learned Noise Schedules
Next, we show the learned noise schedules for the smallest (churn) and the largest (acsincome) datasets. Additionally, we illustrate the fit of single, per type and per feature schedules to the respective losses.
@@
Appendix P Qualitative Comparisons
@@
Appendix Q Visualizations of Captured Correlations
@@
Appendix R Detailed Results
CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes. SMOTE takes too long to sample datasets of a sufficient size for acsincome and covertype (see Table 29). For those models, the performance metrics on these datasets are therefore not reported. They are assigned a rank of 10 in Table 1 and are not taken into account when forming the average metrics reported in Table 11.
Table 11: Model evaluation results averaged over 11 datasets (skipping a dataset if the model was not trainable on it, which due to extensive sampling times for SMOTE includes two of the most complex datasets, acsincome and covertype) for seven benchmark models and for CDTD with three different noise schedules. Per performance metric, bold indicates the best, underline the second best result.
@@
RMSE (abs. diff.; $\downarrow$) 0.083 0.094 0.674 0.947 0.486 0.173 0.313 0.084 0.101 0.110
F1 (abs. diff.; $\downarrow$) 0.007 0.053 0.130 0.074 0.015 0.044 0.099 0.025 0.020 0.025
AUC (abs. diff.; $\downarrow$) 0.008 0.020 0.080 0.065 0.009 0.027 0.059 0.018 0.016 0.022
L2 distance of corr.  ($\downarrow$) 0.866 1.321 2.187 2.745 3.786 1.200 2.025 0.782 0.756 0.990
Detection score ($\downarrow$) 0.661 0.934 0.986 0.976 0.769 0.936 0.877 0.796 0.768 0.783
JSD ($\downarrow$) 0.055 0.011 0.114 0.152 0.051 0.038 0.048 0.015 0.016 0.018
WD ($\downarrow$) 0.004 0.011 0.023 0.025 0.061 0.022 0.016 0.010 0.007 0.009
DCR (abs. diff. to test; $\downarrow$) 1.278 1.588 3.336 1.621 0.568 1.000 2.593 0.796 0.806 0.758
@@
acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.524_{\pm 0.010}$ $0.141_{\pm 0.003}$ $0.129_{\pm 0.003}$ $0.119_{\pm 0.002}$
adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.156_{\pm 0.006}$ $0.493_{\pm 0.009}$ $0.449_{\pm 0.011}$ $0.170_{\pm 0.007}$ $0.125_{\pm 0.009}$ $0.128_{\pm 0.010}$
bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.898_{\pm 0.025}$ $0.499_{\pm 0.021}$ $0.677_{\pm 0.015}$ $0.323_{\pm 0.008}$ $0.266_{\pm 0.011}$ $0.256_{\pm 0.015}$
beijing $0.081_{\pm 0.007}$ $0.133_{\pm 0.006}$ $1.445_{\pm 0.009}$ $1.642_{\pm 0.015}$ $1.133_{\pm 0.035}$ $0.363_{\pm 0.015}$ $0.096_{\pm 0.008}$ $0.075_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.071_{\pm 0.004}$
churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.327_{\pm 0.044}$ $0.746_{\pm 0.062}$ $0.509_{\pm 0.053}$ $0.302_{\pm 0.041}$ $0.289_{\pm 0.043}$ $0.282_{\pm 0.044}$
covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.044_{\pm 0.001}$ $1.029_{\pm 0.032}$ $3.958_{\pm 0.243}$ $2.359_{\pm 0.011}$ $2.275_{\pm 0.009}$ $2.710_{\pm 0.009}$
default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $3.408_{\pm 0.105}$ $1.672_{\pm 0.061}$ $1.121_{\pm 0.042}$ $0.627_{\pm 0.068}$ $0.652_{\pm 0.102}$ $0.737_{\pm 0.033}$
diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.381_{\pm 0.026}$ $1.201_{\pm 0.020}$ $0.803_{\pm 0.032}$ $1.345_{\pm 0.016}$
lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.675_{\pm 0.015}$ $6.701_{\pm 0.034}$ $1.042_{\pm 0.075}$ $1.189_{\pm 0.040}$ $1.363_{\pm 0.097}$
news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $15.985_{\pm 0.081}$ $4.874_{\pm 0.148}$ $4.990_{\pm 0.024}$ $1.925_{\pm 0.527}$ $2.035_{\pm 0.475}$ $3.395_{\pm 0.950}$
nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.447_{\pm 0.031}$ $0.609_{\pm 0.032}$ $0.867_{\pm 0.046}$ $0.433_{\pm 0.025}$ $0.478_{\pm 0.083}$ $0.481_{\pm 0.058}$
@@
acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.000}$ $0.022_{\pm 0.001}$
adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.034_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.020_{\pm 0.001}$ $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.016_{\pm 0.000}$
bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.029_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.054_{\pm 0.001}$ $0.010_{\pm 0.000}$ $0.009_{\pm 0.001}$ $0.012_{\pm 0.001}$
beijing $0.006_{\pm 0.002}$ $0.005_{\pm 0.002}$ $0.147_{\pm 0.003}$ $0.257_{\pm 0.001}$ $0.035_{\pm 0.003}$ $0.018_{\pm 0.004}$ $0.007_{\pm 0.002}$ $0.003_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.005_{\pm 0.001}$
churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.014_{\pm 0.004}$ $0.043_{\pm 0.001}$ $0.017_{\pm 0.002}$ $0.012_{\pm 0.003}$ $0.012_{\pm 0.002}$ $0.011_{\pm 0.002}$
covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.049_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.011_{\pm 0.000}$
default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.027_{\pm 0.002}$ $0.073_{\pm 0.002}$ $0.082_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.095_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.023_{\pm 0.000}$ $0.026_{\pm 0.000}$
lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.243_{\pm 0.002}$ $0.114_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.064_{\pm 0.002}$
news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.046_{\pm 0.000}$ $0.012_{\pm 0.001}$ $0.016_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.028_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.026_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.013_{\pm 0.003}$

### v4

@@ Appendix I Tuning of the Detection Model
These days, inspired from diffusion models in the image and video domains, much work relies on the idea of latent diffusion. Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space alleviates those.
@@
Latent diffusion models first encode the data in a latent space. The diffusion model itself is then trained in that latent space instead of directly on the features. Hence, the performance of the diffusion model directly depends on a second model, with a separate training procedure. TabSyn uses a VAE model to encode mixed-type data into a common continuous space that is not lower-dimensional, so as to minimize reconstruction errors. Any reconstruction errors caused by the VAE reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model. This suggests that we would want to train a highly capable encoder/decoder, which adds additional training costs. Figure 3 shows that latent diffusion is not necessarily more efficient in the tabular data domain. In particular, if the latent space is not lower-dimensional or the encoder/decoder is very complex, then sampling speed is not improved.
@@
We further hypothesize that much tabular data, due to the lack of redundancy and spatial or sequential correlation, is difficult to summarize efficiently in a joint latent space. Hence, compared to other domains, larger VAEs and higher-dimensional latent spaces are required, increasing the training time. Also, there is the risk of the VAE not picking up on subtle correlations within the data or distorting existing correlations by mapping into the latent space. Any correlations not properly encoded cannot be learned by the diffusion model. Since we optimize the VAE on an average loss, its reconstruction and encoding performance of, for instance, minority classes or extreme values in long-tailed distributions is likely lacking. This makes the job of the diffusion model more difficult, if not sometimes impossible.
@@
Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix B and C). This is a crucial part of modeling mixed-type data. Using a VAE to define a diffusion process in latent space only shifts the necessity for homogenization to the VAE training process. Not balancing different feature- or data-types and their losses induces implicit importance weights. Thus, the VAE may sacrifice the reconstruction quality of some features in favor of others (Kendall et al., 2018; Ma et al., 2020).
@@
To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine feature-specific sample quality metrics and those that directly benefit from all features being generated well. Our results in Table N show that latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix H). In particular, the attained maximum univariate metrics as well as the detection score and the L2 distance of the correlation matrices indicate that TabSyn has issues modeling all features and their correlations sufficiently well. This supports our argument that a homogenization of data types is important and that an encoding in latent space may complicate the learning of joint data characteristics.
@@
Table 9: A comparison of the CDTD model to latent diffusion (TabSyn). We average each metric over five sampling seeds and as a robust measure report the median over the 11 datasets. Abs. diff. in corr. matrices refers to the absolute differences in the correlation matrices between ground truth and synthetic data. The max, min and mean are taken across features.
{NiceTabular}
[…]
@@
TabSyn 0.859 0.919 0.004 0.044 0.141 0.002 0.012 0.025 0.000 0.261
CDTD (per type) 0.701 0.444 0.002 0.011 0.025 0.001 0.006 0.010 0.000 0.088
Improv. over TabSyn 18.4% 51.7% 50.0% 75.0% 82.3% 50.0% 50.0% 60.0% - 66.3%
Appendix O Comparison to Related Work
Table O summarizes our comparison of CDTD to the diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi. Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix N). TabSyn is the only other model besides CDTD that avoids one-hot encoding categorical features by using embeddings. This improves the scalability to a higher number of categories without blowing up the input dimensions. Although both models utilize embeddings, TabSyn’s generative capabilities are more constrained by jointly encoding all features in a latent space. It should also be noted that TabSyn makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
CDTD is the first model to utilize adaptive and type- or feature-specific noise schedules to model tabular data. Further, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix B and C). No other model attempts balancing the different features types. This is problematic as it suggests that other models may suffer from feature-specific implicit importance weights that impact both training and generation. Hence, the sample quality of some features may be unintentionally sacrificed in favor of increasing the sample quality of other features (Kendall et al., 2018; Ma et al., 2020). Note that this also applies to TabSyn: Even though their diffusion model avoids this issue by relying on a single type of loss due to the continuous latent space, the VAE training process does not account for any balancing issues between the two data types. Hence, the balancing issue is not eliminated but got only shifted to the VAE.
@@
{NiceTabular}
[…]
CoDi ✓
TabDDPM ✓
TabSyn ✓ $\ast$
CDTD (ours) ✓ ✓ ✓ ✓ ✓ ✓
Appendix P Examples of Learned Noise Schedules
Next, we show the learned noise schedules for the largest (acsincome) and the smallest (churn) datasets. Additionally, we illustrate the fit of single, per type and per feature schedules to the respective diffusion losses they were trained to fit.
@@
Appendix Q Qualitative Comparisons
@@
Appendix R Visualizations of Captured Correlations
@@
Appendix S Detailed Results
CoDi is prohibitively expensive to train on lending and diabetes and TabDDPM produces NaNs for acsincome and diabetes. SMOTE takes too long to sample datasets of a sufficient size for acsincome and covertype (see Table 29). For those models, the performance metrics on these datasets are therefore not reported. They are assigned a rank of 10 in Table 1 and the worst metric-specific performance across all models before computing the average metrics reported in Table 11.
Table 11: Model evaluation results averaged over 11 datasets (if a model was not trainable on a given dataset, we assign the maximum, i.e., worst, value over all models for that dataset to this model) for seven benchmark models and for CDTD with three different noise schedules. Per performance metric, bold indicates the best, underline the second best result.
@@
RMSE (abs. diff.; $\downarrow$) 0.366 0.091 0.635 0.868 0.674 0.278 0.309 0.086 0.091 0.102
F1 (abs. diff.; $\downarrow$) 0.068 0.053 0.130 0.074 0.020 0.048 0.111 0.020 0.015 0.014
AUC (abs. diff.; $\downarrow$) 0.043 0.020 0.080 0.065 0.023 0.039 0.066 0.015 0.014 0.015
L2 distance of corr.  ($\downarrow$) 1.287 1.321 2.190 2.707 3.001 2.383 2.097 0.792 0.684 0.920
Detection score ($\downarrow$) 0.869 0.933 0.986 0.977 0.790 0.947 0.858 0.761 0.739 0.770
JSD ($\downarrow$) 0.077 0.011 0.101 0.135 0.086 0.073 0.052 0.015 0.016 0.018
WD ($\downarrow$) 0.011 0.011 0.024 0.023 0.050 0.059 0.017 0.010 0.008 0.009
DCR (abs. diff. to test; $\downarrow$) 1.813 1.588 3.317 1.602 2.061 2.838 2.648 0.927 0.867 0.760
@@
acsincome - $0.242_{\pm 0.002}$ $1.696_{\pm 0.008}$ $1.136_{\pm 0.004}$ $0.517_{\pm 0.006}$ $0.560_{\pm 0.005}$ $0.134_{\pm 0.002}$ $0.135_{\pm 0.002}$ $0.123_{\pm 0.002}$
adult $0.414_{\pm 0.016}$ $0.576_{\pm 0.006}$ $1.858_{\pm 0.010}$ $0.735_{\pm 0.012}$ $0.160_{\pm 0.012}$ $0.493_{\pm 0.009}$ $0.514_{\pm 0.013}$ $0.175_{\pm 0.007}$ $0.125_{\pm 0.005}$ $0.123_{\pm 0.011}$
bank $0.404_{\pm 0.015}$ $0.819_{\pm 0.024}$ $0.947_{\pm 0.019}$ $2.758_{\pm 0.049}$ $0.529_{\pm 0.050}$ $0.499_{\pm 0.021}$ $0.759_{\pm 0.013}$ $0.333_{\pm 0.018}$ $0.231_{\pm 0.009}$ $0.288_{\pm 0.013}$
beijing $0.081_{\pm 0.007}$ $0.128_{\pm 0.009}$ $1.470_{\pm 0.007}$ $1.226_{\pm 0.008}$ $0.368_{\pm 0.085}$ $0.373_{\pm 0.008}$ $0.086_{\pm 0.008}$ $0.073_{\pm 0.009}$ $0.073_{\pm 0.007}$ $0.071_{\pm 0.009}$
churn $0.264_{\pm 0.036}$ $0.635_{\pm 0.026}$ $1.355_{\pm 0.043}$ $1.301_{\pm 0.041}$ $0.273_{\pm 0.069}$ $0.746_{\pm 0.062}$ $0.613_{\pm 0.022}$ $0.269_{\pm 0.028}$ $0.255_{\pm 0.021}$ $0.271_{\pm 0.040}$
covertype - $1.192_{\pm 0.017}$ $3.685_{\pm 0.005}$ $4.668_{\pm 0.003}$ $1.124_{\pm 0.183}$ $1.029_{\pm 0.032}$ $3.749_{\pm 0.181}$ $1.970_{\pm 0.010}$ $1.357_{\pm 0.155}$ $1.972_{\pm 0.011}$
default $0.709_{\pm 0.048}$ $1.228_{\pm 0.021}$ $2.697_{\pm 0.021}$ $1.564_{\pm 0.029}$ $0.685_{\pm 0.131}$ $1.672_{\pm 0.061}$ $1.125_{\pm 0.049}$ $0.724_{\pm 0.116}$ $0.641_{\pm 0.130}$ $0.681_{\pm 0.037}$
diabetes $2.355_{\pm 0.026}$ $1.189_{\pm 0.004}$ $1.654_{\pm 0.008}$ $5.351_{\pm 0.095}$ $2.796_{\pm 0.066}$ $1.381_{\pm 0.016}$ $1.213_{\pm 0.029}$ $1.351_{\pm 0.044}$
lending $1.321_{\pm 0.063}$ $3.473_{\pm 0.057}$ $2.420_{\pm 0.016}$ $5.895_{\pm 0.026}$ $10.046_{\pm 0.007}$ $6.792_{\pm 0.034}$ $1.148_{\pm 0.087}$ $1.239_{\pm 0.090}$ $1.351_{\pm 0.050}$
news $1.684_{\pm 1.466}$ $4.333_{\pm 0.128}$ $4.641_{\pm 0.028}$ $4.612_{\pm 0.016}$ $12.356_{\pm 0.097}$ $4.874_{\pm 0.148}$ $5.153_{\pm 0.014}$ $2.050_{\pm 0.594}$ $1.811_{\pm 0.295}$ $3.446_{\pm 1.111}$
nmes $0.565_{\pm 0.047}$ $0.717_{\pm 0.054}$ $1.663_{\pm 0.035}$ $0.532_{\pm 0.030}$ $0.426_{\pm 0.041}$ $0.609_{\pm 0.032}$ $0.919_{\pm 0.067}$ $0.454_{\pm 0.043}$ $0.444_{\pm 0.075}$ $0.445_{\pm 0.071}$
@@
acsincome - $0.013_{\pm 0.001}$ $0.256_{\pm 0.000}$ $0.309_{\pm 0.000}$ $0.076_{\pm 0.001}$ $0.052_{\pm 0.001}$ $0.021_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.019_{\pm 0.000}$
adult $0.064_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.112_{\pm 0.001}$ $0.113_{\pm 0.001}$ $0.035_{\pm 0.001}$ $0.045_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.011_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
bank $0.039_{\pm 0.001}$ $0.004_{\pm 0.000}$ $0.086_{\pm 0.001}$ $0.191_{\pm 0.001}$ $0.021_{\pm 0.001}$ $0.038_{\pm 0.001}$ $0.063_{\pm 0.001}$ $0.011_{\pm 0.000}$ $0.011_{\pm 0.001}$ $0.015_{\pm 0.001}$
beijing $0.006_{\pm 0.002}$ $0.004_{\pm 0.001}$ $0.005_{\pm 0.002}$ $0.074_{\pm 0.003}$ $0.024_{\pm 0.002}$ $0.011_{\pm 0.003}$ $0.011_{\pm 0.003}$ $0.006_{\pm 0.002}$ $0.006_{\pm 0.001}$ $0.007_{\pm 0.002}$
churn $0.012_{\pm 0.004}$ $0.011_{\pm 0.004}$ $0.095_{\pm 0.003}$ $0.048_{\pm 0.004}$ $0.015_{\pm 0.006}$ $0.043_{\pm 0.001}$ $0.031_{\pm 0.002}$ $0.011_{\pm 0.002}$ $0.010_{\pm 0.002}$ $0.012_{\pm 0.003}$
covertype - $0.002_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.043_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.008_{\pm 0.000}$ $0.044_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.007_{\pm 0.000}$
default $0.042_{\pm 0.001}$ $0.008_{\pm 0.001}$ $0.194_{\pm 0.001}$ $0.177_{\pm 0.001}$ $0.028_{\pm 0.001}$ $0.073_{\pm 0.002}$ $0.093_{\pm 0.001}$ $0.012_{\pm 0.002}$ $0.016_{\pm 0.001}$ $0.017_{\pm 0.001}$
diabetes $0.067_{\pm 0.000}$ $0.009_{\pm 0.000}$ $0.093_{\pm 0.000}$ $0.187_{\pm 0.000}$ $0.098_{\pm 0.000}$ $0.024_{\pm 0.000}$ $0.025_{\pm 0.000}$ $0.031_{\pm 0.000}$
lending $0.143_{\pm 0.001}$ $0.049_{\pm 0.002}$ $0.092_{\pm 0.001}$ $0.188_{\pm 0.001}$ $0.287_{\pm 0.002}$ $0.119_{\pm 0.001}$ $0.056_{\pm 0.001}$ $0.057_{\pm 0.001}$ $0.065_{\pm 0.002}$
news $0.063_{\pm 0.001}$ $0.002_{\pm 0.001}$ $0.022_{\pm 0.001}$ $0.128_{\pm 0.001}$ $0.017_{\pm 0.001}$ $0.012_{\pm 0.001}$ $0.017_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.060_{\pm 0.001}$ $0.008_{\pm 0.002}$ $0.117_{\pm 0.002}$ $0.029_{\pm 0.003}$ $0.025_{\pm 0.004}$ $0.027_{\pm 0.003}$ $0.019_{\pm 0.001}$ $0.009_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.012_{\pm 0.004}$